Sixty-five randomly selected car salespersons were asked the number of cars they generally sell in one week. Fourteen people answered that they generally sell four cars; nineteen generally sell five cars; twelve generally sell six cars; nine generally sell seven cars; eleven generally sell eight cars. Use this information to calculate the following value. (Enter an exact number as an integer, fraction, or decimal.)

10th percentile
Answer #1

cars (x) frequency (f) cf x*f

4 14 14 56
5 19 33 95
6 12 45 72
7    9 54 63
8 11 65 88

here, n = 65

mean = SUM(x*f)/ SUM(f) = 374/65 = 5.75

P10 = (10(n+1)/100)th value of the observation

= (10*66/100)th value of the observation

= (6.6)th value of the observation

= 4 (Answer)

Percentage of data below the given percentile = 10%

please give a thumbs up